# On the Present Status of Inflationary Cosmology

### v1

@@ 10.1 T-models
Here $\phi(x)$ is the inflaton. In the limit $\alpha\to\infty$ the kinetic term becomes the canonical term $-{(\partial_{\mu}\phi)^{2}\over 2}$, as in the original chaotic inflation model (1). One can recover the canonical normalization for any $\alpha$ by solving the equation ${\partial\phi\over 1-{\phi^{2}\over 6\alpha}}=\partial\varphi$. The solution is $\phi=\sqrt{6\alpha}\,\tanh{\varphi\over\sqrt{6\alpha}}$. The full theory, in terms of the canonical variables $\varphi$, becomes a theory with a plateau potential
@@
This simplest model is shown by the prominent vertical yellow band in Fig. 8 of the paper on inflation in the Planck2018 data release Planck:2018jri . In this model, the condition ${V_{0}\over\alpha}\sim 10^{{-10}}$ reads $m\sim 0.6\times 10^{{-5}}$. The small magnitude of this parameter accounts for the small amplitude of perturbations $A_{s}\approx 2.01\times 10^{{-9}}$. No other parameters are required to match the presently existing Planck/BICEP/Keck data in this model. If the inflationary gravitational waves are discovered, their amplitude can be accounted for by the choice of the parameter $\alpha$ in (22).
@@ 10.2 E-models
The second family of $\alpha$-attractors called E-models is given by
@@
As before, one can go to canonical variables, which yields
@@
Figure 10: On the left panel we show the potential of the T-model (20) for $n=1$, $V_{0}=1$, $73\alpha=1,3,$ 7. On the right, there is an E-model potential (26) for the same parameters. The T-model potential with $\alpha=1,n=1$ is very similar to the potential of the Higgs inflation. The E-model potential with $\alpha=1,n=1$ coincides with the potential in the Starobinsky model.
@@ 10.3 Special cases: Poincare disks and $SL(2,\mathbb{Z})$ $\alpha$-attractors
Figure 11: The figure from “Snowmass2021 Cosmic Frontier: Cosmic Microwave Background Measurements White Paper” Chang:2022tzj . It shows the predictions of T-model $\alpha$-attractors with unconstrained values of $\alpha$ (gray area), the predictions for $13\alpha=7,6,5,4,3,2,$ 1 (purple lines), as well as Higgs inflation, $R^{2}$ inflation (red dots). The predictions are for $5747<N_{e}<$ 57.
@@
So far, we presented T- and E-models with a continuous range of values of $\alpha$. One can implement these models in the minimal ${\cal N}=1$ supergravity, where the parameter $3\alpha$ is given by ${2\over 3\alpha}=|{\cal R}_{K}|$. Here $|{\cal R}_{K}|$ is the curvature of Kähler   geometry Ferrara:2013rsa . In the context of the Poincaré hyperbolic disk geometry, representing an Escher disk, $R^{2}_{\rm Esher}=3\alpha$ defines the size of the disk Kallosh:2015zsa . The most interesting B-mode targets in this class of cosmological attractor models are the ones with the discrete values of $13\alpha=7,6,5,4,3,2,$ 1 Ferrara:2016fwe ; Kallosh:2017ced ; Gunaydin:2020ric ; Kallosh:2021vcf . These models of Poincaré disks are inspired by string theory, M-theory, and maximal supergravity. Predictions of these models are shown by purple lines in Fig. 11.
@@
In this context, we should also mention one of the more developed models of string theory inflation, called fibre inflation, see Cicoli:2024bxw and references therein. It was observed in Kallosh:2017wku that from the phenomenological point of view, fibre inflation can be effectively described as an $\alpha$-attractor with $\alpha=2$. This was also confirmed in Cicoli:2024bxw . Therefore, fibre inflation predicts the values of $n_{s}$ and $r$ shown in Fig. 11 by the second from the top purple line corresponding to a Poincaré disk with $63\alpha=$ 6.
@@
Recently, there were some interesting developments in the theory of $SL(2,\mathbb{Z})$ invariant $\alpha$-attractors Casas:2024jbw ; Kallosh:2024ymt ; Kallosh:2024pat ; Kallosh:2024whb ; Carrasco:2025rud . Potentials of these models have an infinitely large inflationary plateau, which has the same shape in the inflaton direction as the potentials of $\alpha$-attractors. But these potentials also have an infinitely large number of ridges, which, at first glance, are too sharp to support inflation, see Fig. 12. However, this apparent sharpness is just an illusion created by hyperbolic geometry. One can show that each of these ridges is physically equivalent to an infinite inflationary plateau. Thus, these potentials describe an infinitely large number of inflationary plateaus.
@@
Figure 12: A typical $SL(2,\mathbb{Z})$ invariant potential in Cartesian coordinates Carrasco:2025rud . The upper part of the figure shows an infinitely large inflationary plateau. The lower part of the figure shows infinitely many sharp ridges.
@@
The intricate structure of the inflationary potential in the models with axion stabilization is shown in Fig. 13 Carrasco:2025rud . The sharp peaks in this figure represent large inflationary plateaus, which are physically equivalent to each other. All stabilized inflationary trajectories are shown by infinitely many axion valleys (dark lines) in the figure. Inflation occurs when the inflaton field rolls down along these valleys.
Thus, the plot of the potentials with axion stabilization in disk variables instantly provides a detailed and complete map of the $SL(2,\mathbb{Z})$ invariant potentials, including a map of all inflationary trajectories available in these theories. After the axion stabilization Carrasco:2025rud , these models have the same (discrete) set of predictions as the $\alpha$-attractor Poincaré disk models shown by the purple lines in Fig. 11.
@@
11 Pre-attractor regime and ACT
@@
At first glance, all of these models must be affected equally since all of them have the same attractor prediction $n_{s}=1-2/N_{e}$. However, as we already mentioned, typically reheating in the Starobinsky model is less efficient than in the Higgs inflation model, which is why $n_{s}$ in the Starobinsky model is expected to be smaller than in the Higgs inflation model, see Fig. 11.
Also, a quick glance at Fig. 14 shows that the predictions of the simplest T-models (yellow lines) and the simplest E-models (red lines) coincide in the attractor limit $r\ll 0.01$, but they are somewhat different in the pre-attrtactor regime at $0.015\lesssim r\lesssim 0.036$ corresponding to the dark purple area favored by ACT. In this range of $r$, the red curve with $N_{e}=60$ practically touches the dark purple area favored by ACT, see Fig. 14. Thus, E-models with $r$ in this range are not incompatible with ACT. For a related discussion of this issue, see also Haque:2025uri .
@@
Figure 15: Attractor trajectories for $V(\phi)\sim f^{2}(\phi)\sim\phi^{8},\phi^{6},\phi^{4},\phi^{3},\phi^{2},\phi,\phi^{2/3}$, left to rught, for $N_{e}=60$. The results of Kallosh:2013tua are superimposed with the ACT results shown in Fig. 4. The predictions of the model $V\sim\phi^{2}$ cross the central part of the dark purple area favored by ACT for $N_{e}=60$, $40.3<\xi<$ 4. One can show that the predictions for the linear potential $V\sim\phi$ go through the central part of this area for $N_{e}=50$.
@@ 12 Quintessential $\alpha$-attractors, reheating and ACT, DES, and DESI
Many of these models, including the model shown in Fig. 16, have a potential that does not have a minimum. The standard mechanism of reheating, where an oscillating inflaton decays into other particles, does not work in these models. There are two basic mechanisms responsible for reheating in quintessential inflation: gravitational reheating Peebles:1998qn and instant preheating Felder:1998vq ; Felder:1999pv . Gravitational reheating is very inefficient; it may take a lot of time Dimopoulos:2017zvq , but one can increase its efficiency by taking into account instant preheating, thus effectively controlling the density of the universe at the end of reheating $\rho_{\rm reh}$, as well as the reheating temperature $T_{\rm reh}$ Akrami:2017cir . Quintessential $\alpha$-attractor models studied in Akrami:2017cir , as well as in the earlier models in Linder:2015qxa ; Dimopoulos:2017zvq ; Dimopoulos:2017tud , had the main purpose of describing evolving dark energy instead of the $\Lambda$CDM model, as in the general case of quintessential inflation Peebles:1998qn . It has been noticed in Dimopoulos:2017zvq ; Akrami:2017cir that the energy density of the inflaton during the lengthy stage of reheating in these models is dominated not by its potential but by its kinetic energy (the stage of kination). Therefore the parameter $w$ in the equation of stage $p=w\rho$ can be very close to $1$,
@@
During the last few years, additional studies of quintessential $\alpha$-attractors were performed, stimulated by recent Dark Energy experiments and hints that Dark Energy might be evolving. The two-shoulder models which were used in new studies Akrami:2020zxw ; Zhumabek:2023wka ; Giare:2024sdl ; Alestas:2024eic were proposed and studied in Akrami:2017cir . They involve the linear exponential dependence of the potential on the geometric field $\phi=\sqrt{6\alpha}\,\tanh{\varphi\over\sqrt{6\alpha}}$.
@@
This model describes dark energy with an evolving equation of state $w$, asymptotically approaching $w=-1$.
@@
An amazing feature of the Exp-II model (36) is that the asymptotic equation of state never reaches the value $-1$ and is given by
@@
The results of the analysis in Akrami:2020zxw ; Zhumabek:2023wka ; Giare:2024sdl ; Alestas:2024eic appear to be consistent with some values of $3\alpha$ related to Poincaré disks and DES and DESI observations at the time when these studies were performed. For example, in Alestas:2024eic it is suggested that the DESI data are compatible with the Exp-I model (35) with $53\alpha=$ 5 or $63\alpha=$ 6.
@@ 13 Inflationary models with $n_{s}>1-2/N_{e}$
This issue was addressed in many publications following the Planck 2018 data release, see e.g. Kallosh:2018zsi ; Kallosh:2019hzo ; Kallosh:2022feu ; Kallosh:2022ggf and references therein. For a while, this could seem to be a purely academic exercise since there was no demonstrated need to look for the models with $n_{s}>1-2/N_{e}$. However, as we already mentioned, the status of all models with $n_{s}=1-2/N_{e}$ shown in Fig. 11 was challenged by the latest ACT data release, which favored $n_{s}\approx 0.974$ Louis:2025tst ; ACT:2025tim . Therefore, in this section, we will study several models with $n_{s}>1-2/N_{e}$.
@@ 13.1 Chaotic inflation with nonminimal coupling to gravity
$n_{s}=0.974,\quad r=0.0092,\ \qquad N_{e}=60\ ,$ (40)
$n_{s}=0.968,\quad r=0.0121,\qquad\ N_{e}=50\ .$ (41)
@@
This explains an increase of $n_{s}$ and $r$ as compared to the predictions of the Starobinsky model and Higgs inflation $n_{s}=1-{2\over N_{e}}$, $r={12\over N_{e}^{2}}$, and with respect to the predictions of $\alpha$-attractors $n_{s}=1-{2\over N_{e}}$, $r={12\alpha\over N_{e}^{2}}$ Kallosh:2013yoa . Numerically, the inflationary predictions of the model (42) for $N_{e}=60$ are compatible with the P-ACT-LB constraints for $40.3<\xi<$ 4 Kallosh:2025rni .
@@
In the next subsection, we will discuss many other models with predictions compatible with the results of the latest ACT data release. They include pole inflation, KKLTI inflation, polynomial $\alpha$-attractors, and hybrid $\alpha$-attractors. These models are described in Kallosh:2019hzo ; Kallosh:2022feu ; Kallosh:2022ggf .
The main advantage of the model (38) is its ultimate simplicity. Just like the Starobinsky model and the model $\tfrac{1}{2}m^{2}\phi^{2}$, the model (38) has only one parameter, $m$, which determines the amplitude of the inflationary perturbations. On the other hand, its generalized version (42) can also provide a good fit to the ACT data while allowing control of the tensor-to-scalar ratio $r$ by changing $\xi$. In this respect, the model (42) is similar to $\alpha$-attractors, where one can control $r$ by changing the parameter $\alpha$ [^2] In this context we should mention that some modified gravity models in the Palatini formalism are also compatible with ACT data, see e.g. Jarv:2017azx ; Antoniadis:2018ywb ; Barman:2023opy ; Dioguardi:2025vci ; Dioguardi:2025mpp . This class of models is very interesting but somewhat less explored. In particular, all models discussed in this paper up to now have supergravity generalizations, but we are unaware of any supergravity versions of inflationary models based on the Palatini formulation.

### v3

@@ 10.1 T-models
Here $\phi(x)$ is the inflaton. In the limit $\alpha\to\infty$, the kinetic term becomes the canonical term $(\partial_{\mu}\phi)^{2}\over 2$}, as in the original chaotic inflation model (1). One can recover the canonical normalization for any $\alpha$ by solving the equation ${\partial\phi\over 1-{\phi^{2}\over 6\alpha}}=\partial\varphi$. The solution is $\phi=\sqrt{6\alpha}\,\tanh{\varphi\over\sqrt{6\alpha}}$. The full theory, in terms of the canonical variables $\varphi$, becomes a theory with a plateau potential
@@
This simplest model is illustrated by the prominent vertical yellow band in Fig. 8 of the paper on inflation in the Planck 2018 data release Planck:2018jri . In this model, the condition ${V_{0}\over\alpha}\sim 10^{{-10}}$ reads $m\sim 0.6\times 10^{{-5}}$. The small magnitude of this parameter accounts for the small amplitude of perturbations $A_{s}\approx 2.01\times 10^{{-9}}$. No other parameters are required to match the presently existing Planck/BICEP/Keck data in this model. If the inflationary gravitational waves are discovered, their amplitude can be accounted for by the choice of the parameter $\alpha$ in (24).
@@ 10.2 E-models
The second family of $\alpha$-attractors, called E-models, is given by
@@
As before, one can go to canonical variables, which yield
@@
Figure 10: On the left panel, we show the potential of the T-model (22) for $n=1$, $V_{0}=1$, $3\alpha=1,3,7$. On the right, there is an E-model potential (28) for the same parameters. The T-model potential with $\alpha=1,n=1$ is very similar to the potential of the Higgs inflation. The E-model potential with $\alpha=1,n=1$ coincides with the potential in the Starobinsky model.
@@ 10.3 Special cases: Poincare disks and $SL(2,\mathbb{Z})$ $\alpha$-attractors
So far, we presented T- and E-models with a continuous range of values of $\alpha$. One can implement these models in the minimal ${\cal N}=1$ supergravity, where the parameter $3\alpha$ is given by ${2\over 3\alpha}=|{\cal R}_{K}|$. Here $|{\cal R}_{K}|$ is the curvature of Kähler   geometry Ferrara:2013rsa . In the context of the Poincaré hyperbolic disk geometry, representing an Escher disk, $R^{2}_{\rm Esher}=3\alpha$ defines the size of the disk Kallosh:2015zsa .
Figure 11: The figure from “Snowmass2021 Cosmic Frontier: Cosmic Microwave Background Measurements White Paper” Chang:2022tzj . It shows the predictions of T-model $\alpha$-attractors with unconstrained values of $\alpha$ (gray area), the predictions for $3\alpha=7,6,5,4,3,2,1$ (purple lines), as well as Higgs inflation, $R^{2}$ inflation (red dots). The predictions are for $47<N_{e}<57$.
The most interesting B-mode targets in this class of cosmological attractor models are the ones with the discrete values of $3\alpha=7,6,5,4,3,2,1$ Ferrara:2016fwe ; Kallosh:2017ced ; Gunaydin:2020ric ; Kallosh:2021vcf . These models of Poincaré disks are inspired by string theory, M-theory, and maximal supergravity. Predictions of these models are shown by purple lines in Fig. 11.
@@
In this context, we should also mention one of the more developed models of string theory inflation, called fibre inflation, see Cicoli:2024bxw and references therein. It was observed in Kallosh:2017wku that from the phenomenological point of view, fibre inflation can be effectively described as an $\alpha$-attractor with $\alpha=2$. This was also confirmed in Cicoli:2024bxw . Therefore, fibre inflation predicts the values of $n_{s}$ and $r$ shown in Fig. 11 by the second from the top purple line corresponding to a Poincaré disk with $3\alpha=6$.
@@
Recently, there were some interesting developments in the theory of $SL(2,\mathbb{Z})$ invariant $\alpha$-attractors Casas:2024jbw ; Kallosh:2024ymt ; Kallosh:2024pat ; Kallosh:2024kgt ; Carrasco:2025rud ; GonzalezQuaglia:2025qem . Potentials of these models have an infinitely large inflationary plateau, which has the same shape in the inflaton direction as the potentials of $\alpha$-attractors. But these potentials also have an infinite number of ridges, which, at first glance, appear too sharp to support inflation (see Fig. 12).
However, this apparent sharpness is just an illusion created by hyperbolic geometry. One can show that each of these ridges is physically equivalent to an infinite inflationary plateau. Thus, these potentials describe an infinitely large number of inflationary plateaus.
@@
Figure 12: A typical $SL(2,\mathbb{Z})$ invariant potential in Cartesian coordinates Carrasco:2025rud . The upper part of the figure shows an infinitely large inflationary plateau. The lower part of the figure shows an infinite number of sharp ridges.
@@
The intricate structure of the inflationary potential in the models with axion stabilization is shown in Fig. 13 Carrasco:2025rud . The sharp peaks in this figure correspond to large inflationary plateaus that are physically equivalent. All stabilized inflationary trajectories are shown by infinitely many axion valleys (dark lines) in the figure. Inflation occurs when the inflaton field rolls down along these valleys.
@@
Thus, the plot of the potentials with axion stabilization in disk variables immediately provides a detailed and complete map of the $SL(2,\mathbb{Z})$- invariant potentials, including a map of all inflationary trajectories available in these theories. After the axion stabilization Carrasco:2025rud , these models have the same (discrete) set of predictions as the $\alpha$-attractor Poincaré disk models shown by the purple lines in Fig. 11.
11 ACT and pre-attractor regime
@@
At first glance, all of these models must be affected equally since all of them have the same attractor prediction $n_{s}=1-2/N_{e}$. However, as we have already mentioned, typically reheating in the Starobinsky model is less efficient than in the Higgs inflation model, which is why $n_{s}$ in the Starobinsky model is expected to be smaller than in the Higgs inflation model, as shown in Fig. 11.
Also, a quick glance at Fig. 14 shows that the predictions of the simplest T-models (yellow lines) and the simplest E-models (red lines) coincide in the attractor limit $r\ll 0.01$, but they are somewhat different in the pre-attractor regime at $0.015\lesssim r\lesssim 0.036$ corresponding to the dark purple area favored by ACT. In this range of $r$, the red curve with $N_{e}=60$ almost touches the dark purple area favored by ACT, as shown in Fig. 14. Thus, E-models with $r$ in this range are not incompatible with ACT. For a related discussion of this issue, see also Haque:2025uri .
Figure 15: Attractor trajectories for $V(\phi)\sim f^{2}(\phi)\sim\phi^{8},\phi^{6},\phi^{4},\phi^{3},\phi^{2},\phi,\phi^{2/3}$, left to rught, for $N_{e}=60$. The results of Kallosh:2013tua are superimposed with the ACT results shown in Fig. 4. The predictions of the model $V\sim\phi^{2}$ cross the central part of the dark purple area favored by ACT for $N_{e}=60$, $0.3<\xi<4$. One can show that the curve corresponding to predictions for the linear potential $V\sim\phi$ crosses the central part of this area for $N_{e}=50$.
@@ 12 Quintessential $\alpha$-attractors, reheating and ACT, DES, and DESI
[…]
@@
Over the last few years, additional studies of quintessential $\alpha$-attractors have been performed, stimulated by recent dark energy experiments and hints that dark energy might be evolving. The two-shoulder models which were used in new studies Akrami:2020zxw ; Zhumabek:2023wka ; Giare:2024sdl ; Alestas:2024eic were proposed and studied in Akrami:2017cir . They involve the linear exponential dependence of the potential on the geometric field $\phi=\sqrt{6\alpha}\,\tanh{\varphi\over\sqrt{6\alpha}}$.
@@
This model describes dark energy using an evolving equation of state $w$, which asymptotically approaches w=1w=-1.
@@
An amazing feature of the Exp-II model (38) is that the asymptotic equation of state never reaches the value 1-1 and is given by
@@
The results of the analysis in Akrami:2020zxw ; Zhumabek:2023wka ; Giare:2024sdl ; Alestas:2024eic appear to be consistent with some values of $3\alpha$ related to Poincaré disks and DES and DESI observations at the time when these studies were performed. For example, in Alestas:2024eic it is suggested that the DESI data are compatible with the Exp-I model (37) with $3\alpha=5$ or $3\alpha=6$.
@@ 13 Inflationary models with $n_{s}>1-2/N_{e}$
This issue was addressed in numerous publications following the Planck 2018 data release; see, e.g., Kallosh:2018zsi ; Kallosh:2019hzo ; Kallosh:2022feu ; Kallosh:2022ggf and references therein. For a while, this could seem to be a purely academic exercise since there was no demonstrated need to look for the models with $n_{s}>1-2/N_{e}$. However, as we already mentioned, the status of all models with $n_{s}=1-2/N_{e}$ shown in Fig. 11 was challenged by the latest ACT data release, which favored $n_{s}\approx 0.974$ Louis:2025tst ; ACT:2025tim . Therefore, in this section, we will study several models with $n_{s}>1-2/N_{e}$.
@@ 13.1 Chaotic inflation with nonminimal coupling to gravity
$n_{s}=0.9733,\quad r=0.0094,\ \qquad N_{e}=60\ ,$ (42)
$n_{s}=0.9679,\quad r=0.0125,\qquad\ N_{e}=50\ .$ (43)
@@
This explains an increase of $n_{s}$ and $r$ as compared to the predictions of the Starobinsky model and Higgs inflation $n_{s}=1-{2\over N_{e}}$, $r={12\over N_{e}^{2}}$, and with respect to the predictions of $\alpha$-attractors $n_{s}=1-{2\over N_{e}}$, $r={12\alpha\over N_{e}^{2}}$ Kallosh:2013yoa . Numerically, the inflationary predictions of the model (44) for $N_{e}=60$ are compatible with the P-ACT-LB constraints for $0.3<\xi<4$ Kallosh:2025rni .
@@
In the next subsection, we will discuss several other models that yield predictions compatible with the results of the latest ACT data release. They include pole inflation, KKLTI inflation, polynomial $\alpha$-attractors, and hybrid $\alpha$-attractors. These models are described in Kallosh:2019hzo ; Kallosh:2022feu ; Kallosh:2022ggf .
The main advantage of the model (40) is its ultimate simplicity. Just like the Starobinsky model and the model $\tfrac{1}{2}m^{2}\phi^{2}$, the model (40) has only one parameter, $m$, which determines the amplitude of the inflationary perturbations.
On the other hand, its generalized version (44) can also provide a good fit to the ACT data while allowing control of the tensor-to-scalar ratio $r$ by changing $\xi$. In this respect, the model (44) is similar to $\alpha$-attractors, where one can control $r$ by changing the parameter $\alpha$. [^2] In this context we should mention that some modified gravity models in the Palatini formalism are also compatible with ACT data, see e.g. Jarv:2017azx ; Antoniadis:2018ywb ; Barman:2023opy ; Dioguardi:2025vci ; Dioguardi:2025mpp . This class of models is very interesting but somewhat less explored. In particular, all models discussed in this paper up to now have supergravity generalizations, but we are unaware of any supergravity versions of inflationary models based on the Palatini formulation.